# FedMKT: Federated Mutual Knowledge Transfer for Large and Small Language Models

### v2

@@ Appendix B Implementation Details of Token Alignment
      Use edit distance as a similarity measure to find the closest token in the Bloom vocabulary to the token in the LLaMa2 vocabulary.
@@
      To align the two tokenization results and determine the optimal matching path, we utilize a dynamic programming algorithm. As an illustration, consider the tokenization outputs from LLaMa2 and Bloom. LLaMa2’s tokenization yields: [’we’, ’util’, ’ize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’]. In contrast, Bloom’s tokenization produces: [’we’, ’utilize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’]. In this instance, seven terms from LLaMa2 align perfectly with those from Bloom, such as "we" and "dynamic". Notably, the LLaMa2 tokens ’util’ and ’ize’ collectively map to the single Bloom token ’utilize’. In scenarios where multiple tokens align to one, like the 2-to-1 case of ’util’ and ’ize’ mapping to ’utilize’, we consider ’utilize’ as a match for ’util’ based on an optimal vocabulary mapping.
@@ D.1 Hyperparameter Settings
SLM Parameters. During training for the four clients, the local epoch E was set to 1. The learning rates $\eta_{\theta}$ were as follows: for "OPT-1.3b", $\eta_{\theta}$=3e-5; for "GPT-2-xlarge", $\eta_{\theta}$=3e-4; for "Bloom-1b1", $\eta_{\theta}$=3e-5; and for "LLaMa-2-1.3b", the same learning rates as for the LLM were used.
@@ D.2 Data Splitting
D.3 Dataset Licenses
For the datasets RTE/WIC/BoolQ/CQA/ARC-E/ARC-C/DialogSum were downloaded from HuggingFaceLhoest et al. (2021) and under Apache License, Version 2.0. For the S-NI dataset, it was from minillmGu et al. (2023) and under MIT License.
D.4 Machine Configuration

### v4

@@ Appendix B Implementation Details of Token Alignment
      Use minimum edit distance(MinED) as a similarity measure to find the closest token in the Bloom vocabulary to the token in the LLaMa2 vocabulary. The recursion function for MinED in dynamic programming is elaborated in Fu et al. (2023).
@@
      To align the two tokenization results and determine the optimal matching path, we utilize a dynamic programming algorithm. As an illustration, consider the tokenization outputs from LLaMa2 and Bloom. LLaMa2’s tokenization yields: [’we’, ’util’, ’ize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’]. In contrast, Bloom’s tokenization produces: [’we’, ’utilize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’]. In this instance, seven terms from LLaMa2 align perfectly with those from Bloom, such as "we" and "dynamic". Notably, the LLaMa2 tokens ’util’ and ’ize’ collectively map to the single Bloom token ’utilize’. In scenarios where multiple tokens align to one, like the 2-to-1 case of ’util’ and ’ize’ mapping to ’utilize’, we consider ’utilize’ as a match for ’util’ based on an optimal vocabulary mapping. Figure 2 illustrates an example of token alignment between LLaMa2 and Bloom.
@@ D.1 Hyperparameter Settings
SLM Parameters. During training for the four clients, the local epoch E was set to 1. The learning rates $\eta_{\theta}$ were as follows: for "OPT-1.3B", $\eta_{\theta}$=3e-5; for "GPT-2-xlarge", $\eta_{\theta}$=3e-4; for "Bloom-1.1B", $\eta_{\theta}$=3e-5; and for "LLaMa-2-1.3B", the same learning rates as for the LLM were used.
@@ D.2 Data Splitting
D.3 Machine Configuration